You forecast a company to have a ROE of 15%, a dividend payout ratio of 30%. The company has a beta of 1.2. The market risk premium is 8% and the risk free rate is 2%. What is company’s intrinsic forward PE ratio based on the formula? If you also know currently the company has a price of $30, and you forecast the company to have a $1 earnings per share. If firms with similar risks in the industry have a PE ratio of 27 with an estimated earnings growth rate of 12%, is the company overvalued or undervalued based on PEG approach?
Answer #1

Intrinsic Forward P/E ratio = r - ROE* (1 - d)

Here, the given values ROE=15%, d=30%,

calculation of 'r' based on the CAPM model as , r = risk free+ beta* market premium

r= 2%+1.2*8%= 11.6%

P/E ratio =  0.3/(0.116-(0.15*0.7)) = 27.27

PEG ratio=P/E earning per share growth rate

PEG = \frac{30/1}{12}= 2.5

As the PEG exceeds one, the stock is overvalued